in the right angle triangle a=6cm, b=8cm Find: 1. the side(c) which I did 10. 2. cos A, 3. tan B, 4. csc A ? I think I got the second and third answer; 36.86 and 53.14. I'm not sure about it.
Answer by edjones(4080)

1. 10 cm (very good)
2. cos(A)=adj/hyp=8/10=4/5
3. tan(B)=opp/adj=8/6=4/3
4. csc(A)=hyp/opp=10/6=5/3
.
The cos^-1 also called the arc cos is the inverse of the cos and tells you how many degrees are in the angle. You found the degrees correctly but that was not the question asked.
.
Ed